Using a mobile phone costs $30 per month, and an additional $0.16 per minute of use.
a) What is the cost for one month if the phone is used for a total of x minutes?
b) What are the smallest and largest numbers of hours you use the phone in a month if the monthly telephone bill is to be between $102 and $126?
1
Expert's answer
2019-10-09T11:06:27-0400

a) The cost for one month if the phone is used for a total of x minutes is: TC = 30 + 0.16x.

b) If the monthly telephone bill is to be between $102 and $126, then the smallest number of hours you use the phone in a month is:

102 = 30 + 0.16x,

x = 72/0.16 = 450 hours,

the largest number of hours you use the phone in a month is:

126 = 30 + 0.16x,

x = 96/0.16 = 600 hours.